OpenStudy (anonymous):

Brandon received a book for his birthday and read 36 pages in 2 hours. Later, he read 54 pages in 3 hours. Does his reading follow a direction variation? Explain

OpenStudy (anonymous):

18 pages per hour in first sitting. 18 pages per hour in second sitting. Seems like there is no variation.

OpenStudy (anonymous):

wouldnt that be a direct variation?

OpenStudy (anonymous):

y = 18x, where x = number of hours and y = number of pages read

OpenStudy (anonymous):

yes funinabox is correct this is direct variation with k = 18

OpenStudy (anonymous):

I am not fresh with terminology. My point is that the rate of reading is constant. :)

OpenStudy (anonymous):

Ratio k: 36/2 = 54/ 3 = 18 Direct variation because the pages increasing when the hour increasing!

OpenStudy (anonymous):

Got it.

OpenStudy (anonymous):

@GT, who got it?

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

Oh, I'm glad I can help the other solver :D

OpenStudy (anonymous):

:)